# Can Internal Software Metrics Predict App Popularity at Launch? Yeas! and Nays!

### v1

@@
Abstract.
Predicting mobile app popularity before release can provide developers with a strategic advantage in a competitive marketplace, yet it remains a challenging problem. This study explores whether internal software metrics, measurable from source code before deployment, can predict an app’s popularity, defined by user ratings (calculated from user reviews) and DownloadsPerYear (yearly downloads). Using a dataset of 446 open-source Android apps from F-Droid, we extract a wide array of features, including system-, class-, and method-level code metrics, code smells, and app metadata. Additional information, such as user reviews, download counts, and uses-permission, was collected from the Google Play Store. We evaluate regression and classification models across three feature sets: a minimal Size-only baseline, a domain-informed Handpicked set, and a Voting set derived via feature selection algorithms. Regression models perform poorly due to skewed data, with low $R^{2}$ scores. However, when reframed as binary classification (Popular vs. Unpopular), results improve significantly. The best model, a Multilayer Perceptron using the Voting set, achieves F1-scores of 0.72. These results suggest that internal code metrics, although limited in their explanatory power, can serve as useful indicators of app popularity. This challenges earlier findings that dismissed internal metrics as predictors of software quality.
@@ Abstract.
ccs: Software and its engineering Empirical software validation
1. Introduction
The rapid evolution of handheld devices, such as smartphones and tablets, along with the development of mobile operating systems, has broadened the scope, complexity, and opportunity in mobile application development. As the user base continues to grow, so does the demand for mobile applications, leading to an increase in the number of vendors (Fu et al., 2013). Mobile application developers primarily rely on prominent platforms such as Google Play and Apple’s App Store to distribute their applications to end users. These platforms provide a centralized marketplace where developers can sell or distribute their products, and users can compare applications, read reviews and ratings, and make informed decisions—essentially serving as ”one-stop shops” (Zhu et al., 2024). Users contribute to the marketplace by providing reviews and ratings that reflect their level of satisfaction (or dissatisfaction) with an application based on personal experience.
@@ 1. Introduction
Previous research (Ali et al., 2017) observed that Android users participate more in rating apps than iOS users, making the Android app stores more suitable for our study. Leveraging data from these stores, this paper investigates whether app popularity can be predicted using only internal software metrics available at the early stages of development—specifically, at an app’s inception. Prior research has demonstrated that internal metrics can reliably estimate external software quality attributes, such as bug-proneness, change-proneness, and the ease of extending functionality. Given that these external factors are strongly linked to app popularity, we pose the following question: can internal software metrics be used to predict key indicators of popularity, such as user ratings and download counts?
@@
We aim to predict app popularity considering two widely adopted popularity indicators: user rating and download count. Accordingly, our study is guided by the following two research questions.
RQ1: Can we predict an app’s rating using internal software metrics?
To address this question, we first applied different regression algorithms to predict an app’s average user rating (on a scale from 1 to 5) based on its internal software metrics. Our results indicate that internal software metrics measured from the initial app version are insufficient for accurately predicting an app’s future rating. We then reformulated the problem as a binary classification task: labeling apps as Popular or Unpopular based on a predefined rating threshold. In this setting, the results are more promising with an F1-score of 0.72. We conclude that source code analysis, when combined with app metadata, is effective in predicting apps by popularity.
RQ2: Can we predict an app’s number of downloads using internal software metrics?
Following a similar approach to RQ1, we first conducted regression analysis on the yearly download count (i.e., downloads per year) rather than the total number of downloads, as the latter can be heavily influenced by an app’s age. Consistent with our findings for user ratings, we observed that accurately predicting the exact number of yearly downloads using internal metrics alone is challenging. However, when we reframed the task as a binary classification problem—i.e., distinguishing between Popular and Unpopular apps based on a predefined number of downloads—the results are much more encouraging with an F1-score of 0.69. In this context, source code analysis, when combined with app metadata, proves effective in predicting an app’s future downloads.
The findings of our study are promising: developers can predict whether their app is likely to be popular even before releasing the first version. This early insight allows developers to take proactive steps to enhance their app’s potential for success, potentially saving time, money, and reputation. Furthermore, app stores could leverage such predictive models to improve app ranking systems and optimize resource allocation. Our results also open new avenues for research, encouraging the development of more refined internal software metrics to further enhance the accuracy of popularity prediction models.
To enable replication and extension, we share our data publicly [^2] https://github.com/SQMLab/AppPopularityPrediction.
@@
2. Related Work and Motivation
@@ 2.1. App Popularity Analysis
Corral et al. (Corral and Fronza, 2015) examined the influence of C&K metrics on mobile app popularity, finding only a weak correlation. However, their study relied on displayed star ratings, an inaccurate approach as ratings can vary by region, app version, and device type (Google, [n. d.]; Apple, [n. d.]). Similarly, Gezici et al. (Gezici et al., 2019b) reported no significant association between internal code quality and external app quality. In contrast, Cruz et al. (Cruz et al., 2019) argued that thorough testing can enhance code quality, which may in turn improve user ratings. The relationship between code quality and app popularity also appears to be app category–dependent, as observed by Sorbo et al. (Di Sorbo et al., 2021). Noei et al. (Noei et al., 2017) found that both device-specific attributes (e.g., memory capacity, display size, battery life) and app-specific characteristics (e.g., UI complexity, code size) can influence user-perceived quality in Android apps. Similarly, Tian et al. (Tian et al., 2015) observed that, compared to low-rated apps, high-rated apps tend to be larger in size, exhibit greater complexity, and benefit from more extensive marketing efforts. Interestingly, Kuttal et al. (Kuttal et al., 2020) highlighted a disconnect between developer and user perspectives: they found a weak correlation between an app’s popularity on GitHub and its ratings in the App Store.
@@
  • The study relied on the dataset of Grano et al. (Grano et al., 2017), which derived code quality metrics from decompiled APKs. However, decompiled code is not equivalent to the original source code, as compilers introduce various optimizations during the APK build process (You et al., 2021). Consequently, critical code metrics—such as code readability—may differ significantly between the original and decompiled versions. More concerningly, decompiled code often differs from the original source code both syntactically and semantically (Harrand et al., 2019), thereby compromising the validity of any analysis based on metrics derived from it.
  • To evaluate model performance, Catolino adopted the 10-fold cross-validation approach. This method is fundamentally flawed in scenarios involving multiple versions of the same entity, which is true for the used dataset. In such scenarios, 10-fold cross-validation approach introduces data leakage: some versions may end up in the training set while others appear in the test set. This leakage leads to overestimated model performance. Pascarella et al. (Pascarella et al., 2020) and Chowdhury et al. (Chowdhury et al., 2024) have examined this issue in detail. Both studies revisited highly accurate method-level bug prediction models and demonstrated that their performance dropped significantly when evaluated using more realistic approaches that prevent overlap between training and test data.
Based on our previous discussion, we identified a subset of research that explored correlations between app attributes and app popularity. However, these findings were often inconsistent, with both positive and negative correlations reported. Moreover, only a limited number of studies attempted to develop predictive models based on app attributes, and those efforts were undermined by methodological flaws that compromised their reliability. Motivated by these gaps, we conduct a new, rigorous study to examine whether app popularity can be predicted solely from internal software metrics. Importantly, we focus exclusively on the first release of each app, ensuring that our model does not rely on historical data, making it particularly valuable for developers seeking to build a strong reputation from the outset.
2.2. Code Metrics
@@
System-level metrics. System-level metrics play a pivotal role in evaluating the long-term maintainability of software systems (Liu et al., 2024). These metrics provide valuable insights into the overall system architecture and its degradation patterns over time (Mo et al., 2016a; Rachow and Riebisch, 2022; Liu et al., 2024). Architectural issues at the system level not only hinder quality but also escalate the cost and complexity of future enhancements. To address these challenges, researchers have proposed various system-level metrics. For instance, Sethi et al. (Sethi et al., 2009) introduced the Independence Level metric, while Mo et al. (Mo et al., 2016a) developed the Decoupling Level metric. Both aim to assess how effectively a software system can be modularized into independent components. Similarly, MacCormack et al. (MacCormack et al., 2006) proposed the Propagation Cost metric, which measures the degree of coupling among source files within a system. Empirical studies have shown that systems with low decoupling scores tend to be more bug-prone and harder to extend with new functionalities (Mo et al., 2016a). Further advancing this area, Cai et al. (Cai and Kazman, 2023) discussed novel architecture-level maintainability metrics and identified architectural design anti-patterns that serve as early warnings of potential structural issues. These anti-patterns assist developers in detecting and resolving architectural flaws before they evolve into critical maintenance challenges.
@@ 2.4. Other Metadata
Feature richness—often inferred from app descriptions or inferred capabilities—has also been explored as a determinant of user appeal. Sarro et al. (Sarro et al., 2018) and Liang et al. (Liang et al., 2017) showed that the number and diversity of claimed features are highly predictive of user ratings. Using NLP and topic modeling, these studies demonstrate that users are more likely to positively rate apps that advertise functional depth and variety, provided these features align with user needs and are well-integrated.
@@ 3. Methodology
Figure 1. Overview of methodology that involves app collection (with source code and ratings), feature extraction, feature selection, and building predictive models.
[…]
Overview of methodology
3.1. App Collection
As we focus exclusively on Android apps, we selected the Google Play Store for collecting popularity metrics—specifically, ratings and number of downloads. Unfortunately, the Play Store does not provide access to source code. Inspired by earlier works in this area (Grano et al., 2017; Zeng et al., 2019), we leveraged the F-Droid repository [^3] https://f-droid.org/en/ last accessed: Apr 18, 2025 to collect the source code of the apps. Since our aim is to predict an app’s popularity at its inception, we required the earliest available version of each app. F-Droid maintains two repositories: the main repository, which contains the latest three versions of each app, and an archive that stores the remaining versions. We therefore collected information—including package name, source code, GitHub repository URL, and so on—from the oldest available version in the archive. This oldest available version may not represent the application’s first version, but it is the closest to the first version that could be reliably collected. Corresponding information, including popularity indicators, genre, advertisement-related information, release date, and other metadata, was retrieved from Google Play. To map apps between F-Droid and Google Play, we used the package name as a common identifier, similar to (Grano et al., 2017). The package name, also known as the application ID, is unique for every application [^4] https://developer.android.com/build/configure-app-module. We further validated this through a manual inspection of 30 applications that share the same package name on both F-Droid and Google Play. For each application, we systematically compared the app icon, description, and promotional images across the two platforms. Although some discrepancies were observed, particularly in cases where not all elements were identical, the similarities were sufficient to unambiguously confirm that the applications are indeed the same. These differences are primarily attributable to the more frequent updates of visual and textual metadata on the Google Play Store. If an app was not available on both platforms, it was excluded from our dataset. After this step, we obtained a collection of 1,347 unique apps.
@@ 3.1. App Collection
3.2. Popularity Indicators
User-provided star ratings and download counts are key indicators of app popularity (Alhejaili and Blustein, 2022; Liptrot et al., 2024). However, the rating displayed on the Play Store is not the actual rating. The displayed rating is not the average of all ratings and also varies by region, device type, and app version (Google, [n. d.]). In fact, Ruiz et al. (Ruiz et al., 2015) reported that app store ratings do not consistently reflect changes in user satisfaction. As such, we decided to calculate the average rating instead of choosing the displayed rating. To do this, we collected all the user reviews for an app. A user review contains data related to star rating, user’s comment, developer’s reply, corresponding app version, etc.
The use of download count as a standalone metric for assessing application popularity is inherently problematic due to its temporal sensitivity. Specifically, consider two applications, each registering 1,000 downloads—if one was released a year ago while the other has been available for a decade, equating their popularity would be misleading. This disparity arises because download-based popularity measures are intrinsically dependent on the age of the application. To empirically substantiate this hypothesis, we compute Kendall’s tau correlation coefficient between application age and download count. The analysis yields a correlation coefficient of 0.31, accompanied by a p-value below the conventional significance threshold of 0.05. These results provide statistically significant evidence that download count is positively correlated with application age. For this reason, we collected the apps’ release data and calculated each app’s age. This way we normalized the age factor and calculated DownloadsPerYear. We also discarded apps that are less than 1 year old to enable robust analysis. After this step, we were left with 503 apps. To provide an overview of the dataset, Figure 2(b) presents the cumulative distribution functions (CDFs) for the size of the applications, measured in lines of code (LOC), and the number of Java files per application.
@@ 3.2. Popularity Indicators
Figure 2. Cumulative distribution functions for (a) the size (LOC) of the applications in log scale and (b) the number of Java files per application in log scale.
[…]
Cumulative distribution functions for (a) the size (LOC) of the applications in log scale and (b) the number of Java files per application in log scale.
3.3. Feature Extraction

### v2

@@
Abstract
Predicting a mobile app’s popularity before its first release can provide developers with a strategic advantage in a competitive marketplace, yet it remains a challenging problem. This study explores the extent to which internal software metrics, measurable from source code before deployment, can predict an app’s popularity (i.e., user ratings and downloads per year) at its inception. For our analysis, we first constructed a rigorously filtered dataset of 446 open-source Java-based Android apps that are available on both F-Droid and Google Play Store. Using app source code from F-Droid, we extracted a wide array of internal metrics, including system-, class-, and method-level code metrics, code smells, and app metadata. Popularity-related information, including reviews and download counts, was collected from the Google Play Store.
We evaluate different regression and classification models across three feature sets: a minimal Size-only baseline, a domain-informed Handpicked set, and a Voting set derived via different feature selection algorithms. Our results show that, for both app ratings and number of downloads, regression models perform poorly due to skewed rating distributions and a highly scattered range of download counts in our dataset. However, when the problem is reframed as a binary classification task (Popular vs. Unpopular), performance improves significantly—the best model, a Multilayer Perceptron, achieves an F1-score of 0.72.
We conclude that, although internal code metrics alone are insufficient for accurately predicting an app’s future popularity, they do exhibit meaningful correlations with it. Thus, our findings challenge prior studies that have entirely dismissed internal metrics as valid indicators of software quality. Instead, our results align with research suggesting that internal code metrics can be valuable when evaluated within the appropriate context—specifically, we found them useful for classification tasks.
Keywords:
@@ Abstract
journal: Empirical Software Engineering
1 Introduction
The rapid evolution of handheld devices, such as smartphones and tablets, along with the development of mobile operating systems, has broadened the scope, complexity, and opportunity in mobile application development. As the user base continues to grow, so does the demand for mobile applications, leading to an increase in the number of vendors (Fu et al., 2013). Mobile application developers primarily rely on prominent platforms such as Google Play and Apple’s App Store to distribute their applications to end users. These platforms provide a centralized marketplace where developers can sell or distribute their products, and users can compare applications, read reviews and ratings, and make informed decisions—essentially serving as ”one-stop shops” (Zhu et al., 2024). Users contribute to the marketplace by providing reviews and ratings that reflect their level of satisfaction or dissatisfaction with an application based on personal experience.
@@ 1 Introduction
Previous research (Ali et al., 2017) observed that Android users participate more in rating apps than iOS users, making the Android app stores more suitable for our study. Leveraging data from these stores, this paper investigates the extent to which app popularity can be predicted using only internal software metrics available at the early stages of development—specifically, at an app’s inception. Prior research has demonstrated that internal metrics can reliably estimate external software quality attributes, such as bug-proneness, change-proneness, and the ease of extending functionality. Given that these external factors are strongly linked to app popularity, we pose the following question: can internal software metrics be used to predict key indicators of popularity, such as user ratings and download counts?
@@
We aim to investigate the potential of internal code metrics in predicting app popularity, using two widely adopted indicators: user rating and download count. Accordingly, our study is guided by the following two research questions.
RQ1: To what extent can internal software metrics predict an app’s user rating?
To address this question, we first applied different regression algorithms to predict an app’s average user rating based on its internal software metrics. Our results indicate that internal software metrics measured from the initial app version perform very poorly in predicting an app’s future rating. However, to observe if internal metrics have any impact on app popularity, we reformulated the problem as a binary classification task—labeling apps as Popular or Unpopular based on a predefined rating threshold. In this setting, the results are more promising with an F1-score of 0.72. This suggests that an app’s internal code and design quality are important factors in eventual app popularity, but not the only important factor. This is not surprising given that app popularity may also depend on aspects such as the types of features offered and how pleasing its UI is, which are unlikely to be correlated with any internal software metrics.
RQ2: To what extent can internal software metrics predict an app’s number of downloads?
Following a similar approach to RQ1, we first conducted regression analysis on the yearly download count (i.e., downloads per year) rather than the total number of downloads, as the latter can be heavily influenced by an app’s age. Consistent with our findings for user ratings, we observed that predicting the exact number of yearly downloads is extremely challenging. However, when we reframed the task as a binary classification problem—i.e., distinguishing between Popular and Unpopular apps based on a predefined number of downloads—the results are more encouraging with an F1-score of 0.69.
In summary, while internal metrics alone are insufficient to accurately predict an app’s future popularity, they do offer meaningful predictive insights (hence the title, moderately mighty). This finding challenges earlier studies (El Emam et al., 2001; Gil and Lalouche, 2017) that dismissed the utility of such metrics entirely. In line with another stream of research (Chowdhury et al., 2022; Landman et al., 2014), we find that context is king when evaluating internal metrics—observing a negative result in one context (e.g., regression, in our case) does not preclude encouraging results in others (e.g., classification). Our results, therefore, reaffirm the importance—albeit in a limited capacity—of monitoring a variety of code-quality indicators. By detecting early warning signs, developers can proactively improve their apps, potentially saving time, resources, and reputation. Similarly, app stores could leverage internal metrics to enhance ranking algorithms and allocate support more effectively.
To enable replication and extension, we share our data publicly. [^2] https://github.com/SQMLab/AppPopularityPrediction
@@
2 Related Work and Motivation
@@ 2.1 App Popularity Analysis
Corral et al. (Corral and Fronza, 2015) examined the influence of C&K metrics on mobile app popularity, finding only a weak correlation. However, their study relied on displayed star ratings, an inaccurate approach as ratings can vary by region, app version, and device type (Google, 2024; Apple, 2024). Similarly, Gezici et al. (Gezici et al., 2019b) reported no significant association between internal code quality and external app quality. In contrast, Cruz et al. (Cruz et al., 2019) argued that thorough testing can enhance code quality, which may in turn improve user ratings. The relationship between code quality and app popularity also appears to be app category–dependent, as observed by Sorbo et al. (Di Sorbo et al., 2021). Noei et al. (Noei et al., 2017) found that both device-specific attributes (e.g., memory capacity, display size, battery life) and app-specific characteristics (e.g., UI complexity, code size) can influence user-perceived quality in Android apps. Similarly, Tian et al. (Tian et al., 2015) observed that, compared to low-rated apps, high-rated apps tend to be larger in size, exhibit greater complexity, and benefit from more extensive marketing efforts. Interestingly, Kuttal et al. (Kuttal et al., 2020) highlighted a disconnect between developer and user perspectives: they found a weak correlation between an app’s popularity on GitHub and its ratings in the App Store.
@@
  • The study relied on the dataset of Grano et al. (Grano et al., 2017), which derived code quality metrics from decompiled APKs. However, decompiled code is not equivalent to the original source code, as compilers introduce various optimizations during the APK build process (You et al., 2021; Zeng et al., 2019). Consequently, critical code metrics—such as code readability—may differ significantly between the original and decompiled versions. More concerningly, decompiled code often differs from the original source code both syntactically and semantically (Harrand et al., 2019), thereby compromising the validity of any analysis based on metrics derived from it.
  • To evaluate model performance, Catolino adopted a 10-fold cross-validation approach. This method is fundamentally flawed in scenarios involving multiple versions of the same entity, which is true for the used dataset. In such scenarios, 10-fold cross-validation approach introduces data leakage: some versions may end up in the training set while others appear in the test set. This leakage leads to overestimated model performance. Pascarella et al. (Pascarella et al., 2020) and Chowdhury et al. (Chowdhury et al., 2024) have examined this issue in detail. Both studies revisited highly accurate method-level bug prediction models and demonstrated that their performance dropped significantly when evaluated using more realistic approaches that prevent overlap between training and test data.
Based on our previous discussion, we identified a subset of research that explored correlations between app attributes and app popularity. However, these findings were often inconsistent, with both positive and negative correlations reported. Moreover, only a limited number of studies attempted to develop predictive models based on app attributes, and those efforts were undermined by methodological flaws that compromised their reliability. Motivated by these gaps, we conduct a new, rigorous study to assess how accurately app popularity can be predicted using only internal software metrics. Importantly, we focus exclusively on the first release of each app, ensuring that our model does not rely on historical data, making it particularly valuable for developers seeking to build a strong reputation from the outset.
2.2 Code Metrics
@@
System-level metrics. System-level metrics play a pivotal role in evaluating the long-term maintainability of software systems (Liu et al., 2024). These metrics provide valuable insights into the overall system architecture and its degradation patterns over time (Mo et al., 2016a; Rachow and Riebisch, 2022; Liu et al., 2024). Architectural issues at the system level not only hinder quality but also escalate the cost and complexity of future enhancements. To address these challenges, researchers have proposed various system-level metrics. For instance, Sethi et al. (Sethi et al., 2009) introduced the Independence Level metric, while Mo et al. (Mo et al., 2016a) developed the Decoupling Level metric. Both aim to assess how effectively a software system can be modularized into independent components. Similarly, MacCormack et al. (MacCormack et al., 2006) proposed the Propagation Cost metric, which measures the degree of coupling among source files within a system. Empirical studies have shown that systems with low decoupling scores tend to be more bug-prone and harder to extend with new functionalities (Mo et al., 2016a). Further advancing this area, Cai et al. (Cai and Kazman, 2023a) discussed novel architecture-level maintainability metrics and identified architectural design anti-patterns that serve as early warnings of potential structural issues. These anti-patterns assist developers in detecting and resolving architectural flaws before they evolve into critical maintenance challenges.
@@ 2.4 Other Metadata
Feature richness, often inferred from app descriptions or inferred capabilities, has also been explored as a determinant of user appeal. Sarro et al. (Sarro et al., 2018) and Liang et al. (Liang et al., 2017) showed that the number and diversity of claimed features are highly predictive of user ratings. Using NLP and topic modeling, these studies demonstrate that users are more likely to positively rate apps that advertise functional depth and variety, provided these features align with user needs and are well-integrated.
@@ 3 Methodology
Figure 1: Overview of methodology that involves app collection (with source code and ratings), feature extraction, feature selection, and building predictive models.
3.1 App Collection
As we focus exclusively on Android apps, we selected the Google Play Store for collecting popularity metrics—specifically, ratings and number of downloads. Unfortunately, the Play Store does not provide access to source code. Inspired by earlier works in this area (Grano et al., 2017; Zeng et al., 2019; Coppola et al., 2019), we leveraged the F-Droid repository [^3] https://f-droid.org/en/ last accessed: Apr 18, 2025 to collect the source code of the apps. Since our aim is to predict an app’s popularity at its inception, we required the earliest available version of each app. F-Droid maintains two repositories: the main repository, which contains the latest three versions of each app, and an archive that stores the remaining versions. We therefore collected information—including package name, source code, GitHub repository URL, and so on—from the oldest available version in the archive. This oldest available version may not represent the application’s first version, but it is the closest to the first version that could be reliably collected. Corresponding information, including popularity indicators, genre, advertisement-related information, release date, and other metadata, was retrieved from Google Play. To map apps between F-Droid and Google Play, we used the package name as a common identifier, similar to (Grano et al., 2017). The package name, also known as the application ID, is unique for every application [^4] https://developer.android.com/build/configure-app-module. We further validated this through a manual inspection of 30 applications that share the same package name on both F-Droid and Google Play. For each application, we systematically compared the app icon, description, and promotional images across the two platforms. Although some discrepancies were observed, particularly in cases where not all elements were identical, the similarities were sufficient to unambiguously confirm that the applications are indeed the same. These differences are primarily attributable to the more frequent updates of visual and textual metadata on the Google Play Store. If an app was not available on both platforms, it was excluded from our dataset. After this step, we obtained a collection of 1,347 unique apps.
@@ 3.1 App Collection
3.2 Popularity Indicators
User-provided star ratings and download counts are key indicators of app popularity (Alhejaili and Blustein, 2022; Liptrot et al., 2024). However, the rating displayed on the Play Store is not the actual rating. The displayed rating is not the average of all ratings and also varies by region, device type, and app version (Google, 2024). In fact, Ruiz et al. (Ruiz et al., 2015) reported that app store ratings do not consistently reflect changes in user satisfaction. As such, we decided to calculate the average rating instead of choosing the displayed rating. To do this, we collected all the user reviews for an app. A user review contains data related to star rating, user’s comment, developer’s reply, corresponding app version, etc.
The use of download count as a standalone metric for assessing application popularity is inherently problematic due to its temporal sensitivity. Specifically, consider two applications, each registering 1,000 downloads—if one was released a year ago while the other has been available for a decade, equating their popularity would be misleading. This disparity arises because download-based popularity measures are intrinsically dependent on the age of the application. To empirically substantiate this hypothesis, we compute Kendall’s tau correlation coefficient between application age and download count. The analysis yields a correlation coefficient of 0.31, accompanied by a p-value below the conventional significance threshold of 0.05. These results provide statistically significant evidence that download count is positively correlated with application age. For this reason, we collected the apps’ release data and calculated each app’s age. This way, we normalized the age factor and calculated DownloadsPerYear. We also discarded apps that are less than 1 year old to enable robust analysis. After this step, we were left with 503 apps. To provide an overview of the dataset, Figure 2 presents the cumulative distribution functions (CDFs) for the size of the applications, measured in lines of code (LOC), and the number of Java files per application.
@@ 3.2 Popularity Indicators
Figure 2: Cumulative distribution functions for (a) the size (LOC) of the applications in log scale and (b) the number of Java files per application in log scale.
3.3 Feature Extraction